# Less Is More: Engineering Challenges of On-Device Small Language Model Integration in a Mobile Application

### 27 Apr 2026 v1

@@
Abstract
@@ 2.2 On-Device Inference Framework Evaluation
Prompt control Full (system + user) Full Predefined tasks only
@@
Device support Any (with CPU fallback) Any (GPU preferred) Pixel + select Samsung
@@
MLC LLM [2] is a community-maintained compiler-based framework originating from CMU research. It offers broad model support via TVM compilation but requires a heavy build toolchain: Rust (for tokenizer cross-compilation), Android NDK, TVM compiler, and per-model compilation steps. Its Android documentation states that emulator testing is unsupported and that known GPU freeze issues exist on Adreno hardware [2]. The model preparation pipeline (clone, compile, package) adds significant friction compared to downloading a pre-converted weight file. For a solo developer on a rapid iteration cycle, this toolchain complexity was a dealbreaker.
AICore / ML Kit GenAI [3] is Android’s system-level foundation model service, providing Gemini Nano through a high-level API. While it offers seamless deployment (model managed by the OS), it is restricted to Gemini Nano, a single model with no custom alternatives. More critically, it provides only predefined task APIs (summarization, proofreading, rewriting) with no raw prompt access, making iterative prompt engineering impossible. Device support is limited to recent Pixel and select Samsung flagships. For a game that needs custom prompts and must run on a broad device range, AICore was unsuitable.
@@ 3.3 Models Under Study
Over the development period, we evaluated nine models. Table˜2 lists all models that appeared in the codebase at any point.
Table 2: Complete inventory of models evaluated during development.
Model Params Size RAM Shipped Outcome
Gemma 4 E2B $\sim$2.6B 2,583 MB $\geq$8 GB Yes Premium tier
Gemma 4 E4B $\sim$4B 3,650 MB $\geq$12 GB No Too large for most devices
Gemma 3 1B $\sim$1B 1,005 MB $\geq$4 GB No Discarded
Qwen3 0.6B $\sim$600M 614 MB $\geq$2 GB Yes Compact tier
Qwen2.5 1.5B $\sim$1.5B 1,600 MB $\geq$4 GB No Discarded
Phi-4 Mini $\sim$3.8B 3,910 MB $\geq$8 GB No Discarded
DeepSeek R1 1.5B $\sim$1.5B 1,830 MB $\geq$4 GB No Discarded
@@
Testing was performed on two physical devices: Google Pixel 7 Pro (12 GB RAM, Tensor G2) and Google Pixel 10 Pro (12 GB RAM, Tensor G5).
Table 3: Benchmark data from physical device testing.
@@ 3.4 Data Collection and Analysis
Each AI-related commit was analyzed and coded for: (a) what failure or limitation triggered the change, (b) what strategy was applied, and (c) whether the strategy resolved the issue or led to further iteration. Commits were then grouped into 14 thematic phases and cross-referenced to identify recurring patterns. The commit messages follow the Conventional Commits specification (feat:, fix:, refactor:, chore:), providing structured metadata that aids classification.
@@ 4.1 Day 1 (April 21): Ambition Meets Reality
Also on Day 1, Qwen3 0.6B was added as a third model option (commit ccb7010): 614 MB, requiring only 2 GB RAM, versus Gemma 3 1B’s 1 GB file and 4 GB RAM requirement.
@@ 5 Findings: Failure Taxonomy
We identified five categories of failures that emerged during the integration of on-device SLMs. Each category is documented with symptoms, root cause, observed frequency, code examples, and commit evidence. Table˜4 provides an overview.
Table 4: Summary of the five failure categories.
@@
F1 Output Format Violations JSON wrapped in code fences, translated keys High Medium
F2 Constraint Violations Wrong word length, wrong language, repetition Very High Critical
F3 Context Quality Degradation Output quality drops after 3–5 generations Medium Medium
@@
F5 Model Selection Instability Model quality varies drastically across families Periodic High
@@ 5.3 F3: Context Quality Degradation
Frequency: Consistently observed after $\sim$3 generations in Qwen3 0.6B sessions. Gemma 4 E2B showed degradation after $\sim$5–7 generations. The degradation was gradual, not sudden: hint quality decreased incrementally before becoming obviously repetitive.
@@ 5 Generate hints only 0 (word gen) Hint quality
This arc also included feature removal:
@@ 7.3 S3: Progressive Prompt Hardening (F2)
The prompt system evolved through four complete rewrites over 5 days, each responding to a specific observed failure. Table˜7 traces this evolution.
@@ 9.1 Design Heuristics for On-Device SLM Integration
H8: Fewer supported models means a smaller bug surface. Each model has unique instruction-following behavior and requires independent prompt tuning. We evaluated 9 models and shipped 2. Each additional model multiplied the testing and prompt engineering burden. Support the minimum viable set of models that covers your device range.
@@ 9.3 Framework Choice as an Enabler
First, full prompt control enabled the four prompt rewrites documented in section˜7. AICore’s predefined task APIs (summarization, proofreading) would have made this iteration impossible. The ability to adjust system prompts, user prompts, temperature, topK, and topP independently was essential. In a development cycle where prompts were rewritten every 1–2 days, any friction in the prompt-to-inference pipeline would have been a bottleneck.
@@
A third benefit emerged during testing: emulator support. LiteRT-LM’s CPU backend works on emulators, enabling rapid iteration without a physical device for basic prompt testing. MLC LLM explicitly documents that emulators are unsupported, and AICore requires specific Pixel hardware. For a solo developer, the ability to test on an emulator during early iteration was a significant productivity advantage.
@@ 9.6 The Gap Between Vision and Reality
Table˜8 shows the gap between what we envisioned on Day 0 and what we actually shipped on Day 5.
@@
The five-day constraint was deliberate. We did not want an open-ended research timeline; we wanted to experience what a developer faces when a product manager says “ship it by Friday.” Every cut in table˜8 was made under that pressure. A longer timeline would not have eliminated the failures, only delayed the reckoning. The core lesson, that SLMs require radical scope reduction, would have emerged regardless.
@@ 9.7 Limitations of the Progressive Simplification Pattern
The progressive simplification arc (table˜5) raises a question: could we have started at the minimal scope and avoided the 4-day iteration? In retrospect, partially. A developer applying heuristics H1, H6, and H7 from the start would begin with curated word lists and hint-only generation, potentially compressing the iteration cycle from 5 days to 1–2 days.
@@ 9.8 Prompt Engineering as Empirical Science
Our prompt evolution (table˜7) confirms that prompt engineering for SLMs is an empirical process, not a design-time activity. Each rewrite was a hypothesis (“if I add CRITICAL, the model will comply with length constraints”) tested against real model outputs. Some hypotheses held up (priority markers improved compliance), others only partially (concrete examples helped but did not eliminate violations), and the overall direction was always toward simplification: fewer instructions, simpler tasks.
@@ 10 Threats to Validity
External validity. Findings are derived from specific models (Gemma 4 E2B, Qwen3 0.6B) running on specific hardware (Pixel 7 Pro, Pixel 10 Pro) via a specific framework (LiteRT-LM). While we expect the failure categories (F1–F5) to generalize across SLMs and frameworks (format violations and constraint violations are inherent to small models, not specific to LiteRT-LM), the specific mitigation strategies may require adaptation for other contexts. The failure frequencies reported are estimates based on development observations, not controlled experiments with statistically significant sample sizes.
@@ 11 Conclusion
11.1 A Call for Edge-First AI
@@
This is not a hypothetical concern. The EU’s General Data Protection Regulation (GDPR) establishes data minimization and purpose limitation as legal requirements [15]. Brazil’s Lei Geral de Proteção de Dados (LGPD) grants users the right to know how their data is processed and to request its deletion [16]. California’s Consumer Privacy Act (CCPA) provides similar protections [17]. These regulations exist because the default behavior of platforms, collecting data broadly and using it permissively, has repeatedly failed to protect users. Cloud AI APIs, where every prompt and response transits through a third party’s infrastructure, are a new vector for the same pattern.

### 06 May 2026 v2

@@
Less Is More: Engineering Challenges of On-Device Small Language Model Integration in a Mobile Application
William Oliveira
Independent Researcher
contact@woliveiras.com
April 2026
@@ 2.2 On-Device Inference Framework Evaluation
Prompt control Full (system + user) Full Task APIs + Prompt API (beta, no sampler config)
@@
Device support Any (with CPU fallback) Any (GPU preferred) Selected flagships (multi-vendor)
@@
MLC LLM [2] is a community-maintained compiler-based framework originating from CMU research. It offers broad model support via TVM compilation but requires a heavy build toolchain: Rust (for tokenizer cross-compilation), Android NDK, TVM compiler, and per-model compilation steps. Its Android documentation explicitly notes that emulator testing is unsupported and documents a known issue on Adreno hardware: models with weight layouts using the _1 suffix trigger a 20–50 second system UI freeze at the prefill stage (first inference initialization); the _0 suffix layout avoids the issue [3]. The model preparation pipeline (clone, compile, package) adds significant friction compared to downloading a pre-converted weight file. For a solo developer on a rapid iteration cycle, this toolchain complexity was a dealbreaker.
AICore / ML Kit GenAI [4] is Android’s system-level foundation model service, providing Gemini Nano through a high-level API. While it offers seamless deployment (model managed by the OS), it is restricted to Gemini Nano, a single model with no custom alternatives. While ML Kit also offers a Prompt API (beta) for free-form text prompts, it exposes no sampler configuration (temperature, topK, topP), making fine-grained inference control impossible. Device support spans selected flagships across multiple vendors (Google, Samsung, Xiaomi, OnePlus, and others), but remains limited to specific high-end models with no CPU fallback for broader coverage (Table 2). For a game requiring fine-grained prompt control and model selection, AICore was unsuitable.
Table 2: ML Kit GenAI Prompt API supported devices by Gemini Nano version, consulted May 5, 2026 [5].
Nano version Vendor Models
nano-v2 Google Pixel 9, Pixel 9 Pro, Pixel 9 Pro XL, Pixel 9 Pro Fold
Honor Magic V5, Magic 7, Magic 7 Pro
iQOO iQOO 13
Motorola Razr 60 Ultra, Razr Ultra 2025
OnePlus OnePlus 13, OnePlus 13s
OPPO Find N5
POCO F7 Ultra, F8 Pro, F8 Ultra, X7 Pro, X8 Pro
realme GT 7 Pro
Samsung Galaxy Z Fold7, Galaxy Z TriFold
Xiaomi 14T Pro, 15, 15T, 15T Pro, 15 Ultra, 17, 17 Ultra, Pad Mini
vivo X200 FE, T4 Ultra
nano-v3 Google Pixel 10, Pixel 10 Pro, Pixel 10 Pro XL, Pixel 10 Pro Fold
Honor Magic 8 Pro
iQOO iQOO 15
Motorola Signature
OnePlus OnePlus 15, OnePlus 15R
OPPO Find X9, Find X9 Pro, Find X8, Find X8 Pro, Reno 14 Pro 5G, Reno 15 Pro 5G, Reno 15 Pro Mini 5G, Reno 15 Pro Max 5G
realme GT 7T
Samsung Galaxy S26, Galaxy S26+, Galaxy S26 Ultra
vivo X200T, X200, X200 Pro, X300, X300 Pro
@@ 3.3 Models Under Study
Over the development period, seven models appeared in the codebase. Table˜3 lists all of them, distinguishing those that were actually downloaded and executed from those that were registered as candidates but never run.
Table 3: Complete inventory of models that appeared in the codebase. Tested indicates the model was downloaded and executed; Shipped indicates it reached the production configuration.
Model Params Size RAM Tested Shipped Outcome
Gemma 4 E2B $\sim$2.6B 2,583 MB $\geq$8 GB Yes Yes Premium tier
Gemma 3 1B $\sim$1B 1,005 MB $\geq$4 GB Yes No Replaced by Qwen3 0.6B (Day 2)
Qwen3 0.6B $\sim$600M 614 MB $\geq$2 GB Yes Yes Compact tier
Gemma 4 E4B $\sim$4B 3,650 MB $\geq$12 GB No No Never downloaded: too large
Phi-4 Mini $\sim$3.8B 3,910 MB $\geq$8 GB No No Never downloaded: too large
Qwen2.5 1.5B $\sim$1.5B 1,600 MB $\geq$4 GB No No Never downloaded: deprioritized
DeepSeek R1 1.5B $\sim$1.5B 1,830 MB $\geq$4 GB No No Never downloaded: deprioritized
Only three models were ever downloaded and executed: Gemma 4 E2B, Gemma 3 1B, and Qwen3 0.6B. The remaining four were registered in a Day 5 consolidation pass (commit da341c6) as evaluation candidates and removed the same day before any inference was attempted (commit 0e56401). The barrier was practical: model files ranged from 1.6 GB to 3.9 GB. At iterative game development pace, waiting 10–30 minutes per download to test a model that might be discarded within hours was not a viable use of time. This is itself an instance of the constraints documented throughout this paper: hardware and network realities shaped the evaluation boundary before a single inference token was generated.
@@
Performance measurements were collected on a Google Pixel 7 Pro (12 GB RAM, Tensor G2). A Pixel 10 Pro (Tensor G5) was available but benchmark data from it was not systematically collected within the 5-day development window. Table˜4 summarizes the measured performance.
Table 4: Benchmark data collected on a Google Pixel 7 Pro (Tensor G2, 12 GB RAM).
@@ 3.4 Data Collection and Analysis
Each AI-related commit was analyzed and coded for: (a) what failure or limitation triggered the change, (b) what strategy was applied, and (c) whether the strategy resolved the issue or led to further iteration. Commits were then grouped into thematic phases and cross-referenced to identify recurring patterns. The commit messages follow the Conventional Commits specification (feat:, fix:, refactor:, chore:), providing structured metadata that aids classification.
@@ 4.1 Day 1 (April 21): Ambition Meets Reality
Also on Day 1, Qwen3 0.6B was added as a third model option (commit ccb7010): 614 MB, requiring only 2 GB RAM, versus Gemma 3 1B’s 1 GB file and 4 GB RAM requirement. A post-game LLM chat feature was also introduced the same day (commit 588aa21), allowing players to converse with the model after completing a puzzle. It would survive until Day 5.
@@ 5 Findings: Failure Taxonomy
We identified five categories of failures that emerged during the integration of on-device SLMs. Each category is documented with symptoms, root cause, observed frequency, code examples, and commit evidence. Table˜5 provides an overview.
Table 5: Summary of the five failure categories. Frequency scale: Occasional << Frequent << Very Frequent << Systematic; Intermittent = condition-dependent.
@@
F1 Output Format Violations JSON wrapped in code fences, translated keys Frequent Medium
F2 Constraint Violations Wrong word length, wrong language, repetition Very Frequent Critical
F3 Context Quality Degradation Output quality drops after 3–5 generations Occasional Medium
@@
F5 Model Selection Instability Model quality varies drastically across families Intermittent High
@@ 5.3 F3: Context Quality Degradation
Frequency: Observed qualitatively across manual testing sessions during development ($N\approx 15$ sessions). Degradation consistently appeared after $\sim$3 generations in Qwen3 0.6B sessions and after $\sim$5–7 generations in Gemma 4 E2B sessions. These are development observations, not controlled measurements. The degradation was gradual, not sudden: hint quality decreased incrementally before becoming obviously repetitive.
@@ 5 Generate hints only 0 (word gen) Hint quality
This arc also included feature removal, summarized in Table˜7:
@@
Answer to RQ2: Five failure categories emerge when running LLM inference locally on Android: (F1) output format violations, where the model wraps valid JSON in markdown or uses non-English keys; (F2) constraint violations, where the model ignores length, language, or uniqueness requirements; (F3) context quality degradation, where output quality drops after repeated generations within a session; (F4) latency incompatibility, where synchronous generation blocks the UI for minutes; and (F5) model selection instability, where behaviour and reliability vary drastically across model families and parameter counts.
@@ 7.3 S3: Progressive Prompt Hardening (F2)
The prompt system evolved through four complete rewrites over 5 days, each responding to a specific observed failure. Table˜8 traces this evolution.
@@ 7.6 S6: Guaranteed Deterministic Fallback (All)
Answer to RQ3: Six strategies effectively mitigate the failure categories: (S1) multi-layer defensive parsing to handle format violations; (S2) contextual retry with failure feedback to improve constraint compliance; (S3) progressive prompt hardening to progressively narrow the model’s output space; (S4) session rotation to reset context degradation before it accumulates; (S5) responsibility reduction to eliminate entire failure classes by shrinking the LLM’s task scope; and (S6) guaranteed deterministic fallback to ensure the feature is always functional regardless of LLM success.
@@ 9.1 Design Heuristics for On-Device SLM Integration
H8: Fewer supported models means a smaller bug surface. Each model has unique instruction-following behavior and requires independent prompt tuning. We registered 7 models, tested 3, and shipped 2. Each additional model multiplied the testing and prompt engineering burden. Support the minimum viable set of models that covers your device range.
@@ 9.3 Framework Choice as an Enabler
First, full prompt control enabled the four prompt rewrites documented in section˜7. Although AICore’s ML Kit now offers a Prompt API (beta) for free-form prompts, it exposes no sampler configuration (temperature, topK, topP), which would have prevented the fine-grained inference control essential for this work. The ability to adjust system prompts, user prompts, temperature, topK, and topP independently was essential. In a development cycle where prompts were rewritten every 1–2 days, any friction in the prompt-to-inference pipeline would have been a bottleneck.
@@
A third benefit emerged during testing: emulator support. LiteRT-LM’s CPU backend works on emulators, enabling rapid iteration without a physical device for basic prompt testing. MLC LLM explicitly documents that emulators are unsupported, and AICore requires specific flagship hardware across select vendors. For a solo developer, the ability to test on an emulator during early iteration was a significant productivity advantage.
@@ 9.6 The Gap Between Vision and Reality
Table˜9 shows the gap between what we envisioned on Day 0 and what we actually shipped on Day 5.
@@
The five-day constraint was deliberate. We did not want an open-ended research timeline; we wanted to experience what a developer faces when a product manager says “ship it by Friday.” Every cut in table˜9 was made under that pressure. A longer timeline would not have eliminated the failures, only delayed the reckoning. The core lesson, that SLMs require radical scope reduction, would have emerged regardless.
@@ 9.7 Limitations of the Progressive Simplification Pattern
The progressive simplification arc (table˜6) raises a question: could we have started at the minimal scope and avoided the 4-day iteration? In retrospect, partially. A developer applying heuristics H1, H6, and H7 from the start would begin with curated word lists and hint-only generation, potentially compressing the iteration cycle from 5 days to 1–2 days.
@@ 9.8 Prompt Engineering as Empirical Science
Our prompt evolution (table˜8) confirms that prompt engineering for SLMs is an empirical process, not a design-time activity. Each rewrite was a hypothesis (“if I add CRITICAL, the model will comply with length constraints”) tested against real model outputs. Some hypotheses held up (priority markers improved compliance), others only partially (concrete examples helped but did not eliminate violations), and the overall direction was always toward simplification: fewer instructions, simpler tasks.
@@ 10 Threats to Validity
External validity. Findings are derived from specific models (Gemma 4 E2B, Qwen3 0.6B) running on specific hardware (Pixel 7 Pro) via a specific framework (LiteRT-LM). A Pixel 10 Pro was available during development but systematic benchmark collection was not performed on it; benchmark data therefore reflects a single-device configuration. While we expect the failure categories (F1–F5) to generalize across SLMs and frameworks (format violations and constraint violations are inherent to small models, not specific to LiteRT-LM), the specific mitigation strategies may require adaptation for other contexts. The failure frequencies reported are estimates based on development observations, not controlled experiments with statistically significant sample sizes.
@@ 11 Conclusion
11.1 Privacy and Economic Implications
@@
Declarations
Funding
This research received no external funding.
Ethical Approval
Not applicable.
Informed Consent
Not applicable.
Author Contributions
William Oliveira: conceptualization, study design, data collection, data analysis, writing (original draft), writing (review and editing).
Data Availability Statement
The complete Git history of the Palabrita project is publicly available at https://github.com/woliveiras/palabrita. A preprint of this manuscript is available on arXiv at https://arxiv.org/abs/2604.24636.
Conflict of Interest
The author declares no conflicts of interest.
Acknowledgments
The author thanks David Chou (Product Manager, Android AI Core, Google) for identifying an inaccuracy in an earlier version of this manuscript regarding the ML Kit Prompt API and its device support, and for permitting this acknowledgment.